please express 9 - 7/8(4) in a fraction
Answer by stanbon(75887)

please express 9 - [7/8(4)] in a fraction
-----
= 9 - 7/32
----
= (9*(32/32)) - 7/32
---
= (288/32) - (7/32)
---
= 381/32
=============
You may have meant the following:
9-(7/8)4
= 9-(7/2)
= (18/2)-(7/2)
= 11/2
= 5 1/2
==========
Cheers,
Stan H.